A) B) 1.4 3.6. D) 2.6. 3 19) You won the lottery and have a number of choices as to how to take the money. Which choice19 yields a greater present value? nooy A) $12,000 a year at the end of each of the next 6 years using a 6% discount rate B) $92,000 (lump sum) 7 years from now using an 8% discountrate 47 yas2 QS53 500 (lump sum) now using a 6% discount rate ts4rtog,444 D) $90,000 (lump sum) 7 years from now using a 6% discount rate 20) When you graduate from college, your mother plans to give you a gift of $50,000 to start you on 20) your way. However, to determine what you learned in business school, your mother presents you with four options on how to receive the gift. Which of the four options presented by your mother will yield the greatest present value to you? A) $25,000 per year for the next 2 years using a 3% discount rate B) A lump sum of $50,000 today C) A hump sum of $50,000 after grad school (2 years) assuming a 3% discount rate lump sum of $50,000 after grad school (2 years) assuming a 5% discount rate 21) Assuming an interest rate of 10%, if you invest a lump sum of $5,000 now, the balance of your 21) investment in 7 years will be closest to B) $9,745 A) $35,000. C) $12,970. D) $23,340. 22) The present value of $1,000,000 received in 13 years, given an interest rate of 3%, is 22) A) $1,469,000. B) $661,000 C) $681,000 D) $10,635,000. 23) On a whim you purchased a scratch-off lottery ticket at the gas station. It must have been your 23 lucky day because you won $1,000,000. Being logical and rational you decide to invest the money at 3% for 10 years until you are ready to start a family. At the end of 10 years, how much will your investment be worth? A) $1,344,000 B) $1,384,000 C) $744,000 D) $11,464,000

I feel like my professor is asking ambiguous questions
Answer #1
19)
A) $ 12000 at end of every year for 6 year 6 % Discount rate
Pv for 6 at rate 6% 4.917
PV of cash infolw = 12000 *4.917 = 59004
B) 92000 Lum sum at end of 7 year 8% Discount rate
pv on end of 7 year for 8% 0.5835
PV of cash inflow = 92000*.5835 = 53682
C) $ 53500 now this is already present value
D) 90000 Lum sum at end of 7 year 6% Discount rate
pv on end of 7 year for 6% 0.665
PV of cash inflow = 90000*.665 = 59850

OPTION D HIGHER PRESENT VALUE

20)
A) $ 25000 per year for next 2 years at rate of 3% discount rate
PV for 2 year at rate of 3% 1.9135
PV OF Cash Flow = 25000 * 1.9135 = 47838
B) $ 50000 now is already present value for today
C) 50000 lum sum end of 2 year discount rate 3 %
pv at the end of 2 year for 3% = 0.943
pv Of cash inflow = 50000*.943 = 47150
D) 50000 lum sum end of 2 year discount rate 5 %
pv at the end of 2 year for 5% = 0.907
pv Of cash inflow = 50000*.907 = 45350
OPTION B HIGHER PRESENT VALUE
21)
Interst 10%
Investment 5000
Balance after 7 year Computing int
IF invest $ 1 at the rate of 10 % after 7 year = 1.949 A=P(1+R)t t DENOTE POWER OF (1+R)
So after 7 year amount will be = 5000*1.949 =9745 ( Answer B is Right)
22)
Present value of 1,000,000 received in 13 years at rate of 3% is
PV at end of 13 years at rate of 3% = 0.681
PV of Cash inflow = 1,000,000 *0.681 = 681000
Answer C IS RIGHT